you can buy 3 hotdogs and 1 coke for $4.55. you can buy 2 hotdogs and 1 coke for $3.30 how much does 1 hotdog cost
Answer by mananth(16946)

x cost of each hot dog
y cost of each coke
Total attended
3 x + 1 y = 4.55 .............1

2.00 x + 1.00 y = 3.30 .............2
Eliminate y
multiply (1)by -1.00
Multiply (2) by 1.00
-3.00 x -1.00 y = -4.55
2.00 x + 1.00 y = 3.30
Add the two equations
-1.00 x = -1.25
/ -1.00
x = 1.25
plug value of x in (1)
3.00 x + 1.00 y = 4.55
3.75 + 1.00 y = 4.55
1.00 y = 4.55 -3.75
1.00 y = 0.80
y = 0.80
$1.25 cost of each hot dog
$0.80 cost of each coke
m.ananth@hotmail.ca